-- EvgueniVlassov - 2015-05-02

HF Beam and source calibrations. Thursday, 27 February 2003, https://indico.cern.ch/event/a03379/session/s4/contribution/s4t10/material/0/0.ppt

HF Study of luminescence of HF fibers by radioactive source. Proposal. Wednesday, Thursday, 27 February 2003 , https://indico.cern.ch/event/a03379/session/s4/contribution/s4t9/material/0/0.ppt

HF Lumi_Proposal, Thursday, 6 March 2003 \\cern.ch\dfs\Users\v\vlassov\Documents\Work\Lumi_propos.ppt

HF Beam and source calibrations. Thursday, 27 February 2003, https://indico.cern.ch/event/a03379/session/s4/contribution/s4t10/material/0/0.ppt

HF Study of luminescence of HF fibers by radioactive source. Proposal. Wednesday, Thursday, 27 February 2003 , https://indico.cern.ch/event/a03379/session/s4/contribution/s4t9/material/0/0.ppt

HF Luminescence of irradiated fibers, Thursday, 5 June 2003 https://indico.cern.ch/event/a031235/session/s4/contribution/s4t13/material/0/0.ppt

HF Source calibration, Tuesday, 16 September 2003 https://indico.cern.ch/event/a032106/session/s12/contribution/s12t10/material/0/0.ppt

HF Source calibration: problems detected so far, Friday, 19 September 2003 https://indico.cern.ch/event/a032106/session/s4/contribution/s4t26/material/0/0.ppt

HF Histogram sourcing error, Thursday, 20 November 2003, https://indico.cern.ch/event/a036378/contribution/s1t0/material/0/0.ppt

HF The problems of HF testbeam setup. Tuesday, 9 December 2003, https://indico.cern.ch/event/a036161/session/s12/contribution/s12t8/material/0/0.ppt

HF Helium in HF PMTs. Thursday, 10 June 2004, https://indico.cern.ch/event/a042169/session/s4/contribution/s4t27/material/0/0.ppt

ECAL CCS software requirements, Lisbon, 7 April 2005, http://indico.cern.ch/event/a051/session/0/contribution/8/material/slides/0.ppt

ECAL CCS Configuration and Conditions database, Lisbon, 7 April, 2005, http://indico.cern.ch/event/a051/session/6/contribution/12/material/slides/0.ppt

ECAL Status at 867 and h4 of DAQ components, Tuesday, 8 November 2005, https://indico.cern.ch/event/a057264/contribution/s1t0/material/0/0.ppt

ECAL Preliminary study of pedestals measurements. Static and dynamic modes. Wednesday, 8 February 2006, https://indico.cern.ch/event/a06968/session/s0/contribution/s0t1/material/0/0.ppt

ECAL CCS firmware status. Friday, 22 February 2008 https://indico.cern.ch/event/29467/contribution/2/material/slides/0.ppt

ECAL FE_timing, 22 May 2008, https://indico.cern.ch/event/34080/contributions/1737632/attachments/669081/919665/FE_timing.ppt

ECAL problematic towers -- DCC/CCS aspects.Tuesday, 14 October 2008 https://indico.cern.ch/event/42793/contribution/1/material/slides/0.ppt

ECAL token rings redundancy. Thursday, 12 February 2009 https://indico.cern.ch/event/51297/contribution/0/material/slides/0.ppt

ECAL How to bypass FENIX TPG bug. Proposal. Thursday, 11 June 2009 https://indico.cern.ch/event/61389/contribution/2/material/slides/0.ppt

ECAL DCU readout with L1As in the token ring. Thursday, 2 July 2009 https://indico.cern.ch/event/62890/contribution/7/material/slides/0.ppt

ECAL DCU measurements during run. Thursday, 23 July 2009, https://indico.cern.ch/event/64946/contribution/2/material/slides/0.ppt

ECAL news from DCU r/o. Thursday, 27 August 2009, https://indico.cern.ch/event/66948/contribution/1/material/slides/0.ppt

ECAL CCS and Front-end. CMS ECAL DAQ and Trigger Readiness Review. Wednesday, 16 September 2009, https://indico.cern.ch/event/63447/session/1/contribution/2/material/slides/0.ppt

ECAL problems with TPLL->QPLL->GOL chain behavior with EB+9/10/16/18 and EB-1/3/10. Thursday, 14 January 2010, https://indico.cern.ch/event/80447/contribution/2/material/slides/0.ppt

ECAL Front End Control workshop. Wednesday, 20 January 2010, https://indico.cern.ch/event/58191/session/5/contribution/14/material/slides/4.ppt

ECAL FENIX and Fine Grain Bit . Friday, 12 February 2010, https://indico.cern.ch/event/84617/contribution/2/material/slides/0.ppt

ECAL i_eta bit. Thursday, 25 February 2010, https://indico.cern.ch/event/85910/contribution/0/material/slides/0.ppt

ECAL Strip Fine Grain Update. Thursday, 11 March 2010, https://indico.cern.ch/event/87683/contribution/1/material/slides/0.ppt

ECAL sFGVB online/emulator comparisons. Thursday, 3 March 2011, https://indico.cern.ch/event/129589/contribution/7/material/slides/0.ppt

ECAL Status of new CCS firmware. Wednesday, 10 August 2011, https://indico.cern.ch/event/150077/contribution/2/material/slides/0.pptx

ECAL Recent token-ring problems. Wednesday, 24 August 2011, https://indico.cern.ch/event/150079/contribution/2/material/slides/0.pptx

ECAL Token ring problems. Wednesday, 14 September 2011, https://indico.cern.ch/event/154939/contribution/5/material/slides/0.pptx

ECAL Tau Jet Et trigger asymmetry. Tuesday, 6 December 2011, https://indico.cern.ch/event/151256/contribution/3/material/slides/0.pptx

ECAL CMS ECAL TR operational issues. CMS electronics week: CMS Token Ring Operation workshop, Monday, 23 January 2012 https://indico.cern.ch/event/164323/contribution/3/material/slides/0.pptx

ECAL CCS FW modifications for 2012. ECAL readiness review Tuesday, 31 January 2012 https://indico.cern.ch/event/169719/session/0/contribution/28/material/slides/0.pptx

ECALCCSSupervisor and JDB status. Wednesday, 8 February 2012, https://indico.cern.ch/event/172638/contribution/2/material/slides/0.pptx

ECAL SEUs with TPLLs reset in 2012. Wednesday, 13 June 2012, https://indico.cern.ch/event/184335/contribution/2/material/slides/0.pptx

ECAL EE-8 puzzle. Wednesday, 22 August 2012, https://indico.cern.ch/event/202440/contribution/3/material/slides/0.pptx

ECAL New problematic channels. Tuesday, 25 September 2012, https://indico.cern.ch/event/197430/contribution/5/material/slides/0.pptx

ECAL Cold start July 17. Wednesday, 31 July 2013, https://indico.cern.ch/event/264597/contribution/0/material/slides/4.pptx

TR-fails-P5. Tuesday, 19 November 2013, https://indico.cern.ch/event/283144/contribution/0/attachments/522080/720118/TR-fails-P5.pdf

ECAL EB+9 token ring 5 problem. Wednesday, 26 March 2014, https://indico.cern.ch/event/309905/contribution/0/material/slides/10.pdf

FE_AGR_results. Wednesday, 16 April 2014, https://indico.cern.ch/event/313130/contribution/1/attachments/599536/825019/FE_AGR_results.pdf

ECAL TCDS timing. Wednesday, 10 September 2014, https://indico.cern.ch/event/339963/contribution/1/material/slides/0.pptx

Timex: ECAL vs TCDS timing. Tuesday, 10 March 2015, https://indico.cern.ch/event/379719/contribution/2/attachments/757572/1039218/ECAL_timing_Mar_10.pdf

OM: FE_April_GR_Results Wednesday, 16 April 2014, https://indico.cern.ch/event/313130/contributions/722956/attachments/599536/825020/FE_AGR_results.pptx

OM: EBp9Tr5-EEp4Tr8. Wednesday, 15 July 2015, https://indico.cern.ch/event/433330/contribution/1/attachments/1125970/1609671/EBp9Tr5-EEp4Tr8.pdf

OM: OOS_run_254979. Wednesday, 26 August 2015, https://indico.cern.ch/event/441094/contribution/0/attachments/1144427/1640568/OOS_run_254979.pdf

OM: SyncLoss. Wednesday, 16 December 2015, https://indico.cern.ch/event/470997/contribution/0/attachments/1205762/1756946/SyncLoss.pdf

OM: Magnet ramps and LV switching off/on, Wednesday, 25 November 2015, https://indico.cern.ch/event/464140/contribution/0/attachments/1194024/1735460/Magnet_ramps_and_LV.pdf

Topic attachments
I Attachment History Action Size Date Who Comment
PowerPointppt Lumi_propos.ppt r1 manage 112.5 K 2015-05-02 - 20:26 EvgueniVlassov HF Luminescence of irradiated fiber sudy proposal
Edit | Attach | Watch | Print version | History: r7 < r6 < r5 < r4 < r3 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r7 - 2021-01-31 - EvgueniVlassov
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback